THE SYMMETRY OF INTERSECTION NUMBERS IN GROUP THEORY

PETER SCOTT

Abstract. For suitable subgroups of a finitely generated group, we define the intersection number of one subgroup with another subgroup and show that this number is symmetric. We also give an interpretation of this number.